Definition, Formula and Physical interpretation of Collinear vector.


1
Expert's answer
2021-11-16T13:29:44-0500

When two or more given vectors lie along the same given line, then they can be considered as collinear vectors. We can consider two parallel vectors as collinear vectors since these two vectors are pointing in exactly the same direction or opposite direction.


Two vectors A and B are collinear if there exists a number n, such that A = n · b. Two vectors are collinear if relations of their coordinates are equal, i.e. x1 / x2 = y1 / y2 = z1 / z2


Physical interpretation:

Collinear vectors are two or more vectors which are parallel to the same line irrespective of their magnitudes and direction.